# A Fuzzy-based Approach to Predict Human Interaction by Functional Near-Infrared Spectroscopy

### v1

@@ IV-A2 Picture Rating
During two experiments, fNIRS signals were recorded from two participants simultaneously at a sampling rate of 7.8125 Hz. This was accomplished using 8 sources and 7 detectors on each participant, resulting in a network of 20 source-detector pair channels (CH) per individual covering the Prefrontal Cortex (PFC). To ensure signal quality, we calculated the scalp coupling index (SCI) for each channel, excluding those with an SCI below 0.5. Subsequently, the data underwent artifact removal and baseline correction, with hemoglobin concentrations computed using the Beer-Lambert law. Finally, all fNIRS data were band-pass filtered between 0.01 Hz and 0.2 Hz.
@@ IV-B Implementation Details
IV-C Evaluation Metrics
@@
As depicted in Table I, we show model performance across two dataset arrangements, each with the channel-first and time-first structures, to investigate the effect of input structuring on the classification outcomes. The dimensional reversal aims to explore the robustness of the model against variations in the temporal-spatial data arrangement and explore the explanation capacity of the Fuzzy Attention Layer.
Given that this study pioneers the classification of handholding conditions using fNIRS data, we establish a baseline with bidirectional LSTM and Transformer models, which are state-of-the-art architectures known for their efficacy in sequence modeling tasks. The Fuzzy Attention Layer-based model achieves the Top 1 in PR AUC metrics in four experiments.
We observe that centering in the time dimension significantly enhances the model’s ability to interpret fNIRS data, which adheres to the Haemodynamic Response Function (HRF) [63]. The typical HRF is shown in Figure 4. This adherence to the HRF is critical as HRF-like data patterns exhibit stable temporal characteristics that are optimally suited for analysis by computational models. Particularly, implementing a fuzzy attention layer within our model shows substantial improvements in performance, as it can more effectively learn and adapt to these HRF-like patterns. Figure 1C demonstrates the HRF-like pattern that the center has learned to recognize. This capability underscores the model’s enhanced interpretability and robustness when handling time-first fNIRS data.
@@
IV-D Ablation Study
IV-D1 Exploring the Impact of Fuzzy Attention Layer Replacements within Transformer Encoders
@@
IV-D2 Investigating the Rule Count for Model Optimization
The efficacy of the model is analyzed with respect to the number of rules employed within the computational framework. Figure 5 delineates the relationship between the PR AUC and the rule count, indicating a nuanced dependency on the number of rules. Initially, the PR AUC exhibits an enhancement as the number of rules increases, peaking at approximately ten rules. After that, the model’s performance demonstrates a declining trajectory post the apex, characterized by fluctuations in the PR AUC. This downturn signals a saturation point, beyond which further additions of rules contribute to overfitting. The increase of an excessive number of rules escalates the model’s complexity and parameter count, which, although aimed at refining precision on training data, paradoxically impairs its generalization capabilities on novel, unseen datasets. Consequently, the model exhibits diminished efficacy, underscoring the critical balance required in the rule count to maintain optimal performance without succumbing to overfitting.
@@ IV-D2 Investigating the Rule Count for Model Optimization
IV-D3 Impact of Model Depth on Encoder Performance
@@ V-A Fuzzy Set and Feature Distribution
The fuzzy set distributions of each fuzzy rule, as depicted in Figure 6, illustrate the temporal dynamics and channel specificity of fuzzy membership functions over multiple experimental conditions. The channels 18, 20, and 29 graphs highlight the variability in attention across different time intervals (3s, 5s, 7s, 10s). Each colored line represents a distinct rule derived from fuzzy logic that interprets the fNIRS data, which is overlaid as blue histograms.
@@
This analysis underscores the utility of fuzzy logic-based rules in dissecting the complex temporal patterns of brain activity recorded through fNIRS, offering insights into the nuanced mechanisms of attention distribution during cognitive tasks.
@@ V-B Sample-wised Interpretability Analysis
To provide an intuitive illustration of how the Fuzzy Attention Layer identifies the handholding condition from fNIRS data, we present a demo sample from the top-performing model, where the subject holds a hand.
@@ V-C1 Channel-wise Analysis
We selected the model demonstrating the highest performance for the Picture Recognition dataset, which included 10 rules. We chose this top-performing model and used a t-test to analyze the differences in firing strength. The t-values describe the differences between handholding and non-handholding in the PFC, as shown in Figure 2A. The results indicate how the rules contribute to classification. For example, rules #1, #2, #3, #5, #9, and #10 involve HbO in the right lPFC.
@@ V-C2 Time-wise Analysis
Similarly, in the Picture Recognition dataset, we employed a t-test to evaluate the differences in firing strength between the labels. The t-values illustrating these differences are depicted in Figure 2B. The results indicate that the crucial time domain is approximately 5 seconds after stimuli presentation in rules #5, #6, #7, and #8, which is highly related to the HRF. Rule #5, in particular, shows the most significant difference between handholding and non-handholding conditions(p < 0.05). Figure 2C displays the center prototype in PFC for both oxygenated (HbO) and deoxygenated hemoglobin (HbR), and we also observe the left prefrontal cortex (lPFC)’s contribution in the fuzzy centers noted in the Channel-wise model observation.
@@ V-C3 IBS Analysis
The results indicate that both correlation and cosine similarity metrics showed significantly higher IBS in the handholding condition compared to the non-handholding condition ($\textit{p}<0.05$).
@@ VI Conclusion
In this work, we introduce a new attention layer, the Fuzzy Attention Layer, to identify interpretable patterns of neural activity, which are crucial for detailed analysis and understanding of the decision-making process of the neural network for individual cases. This method enhances the correlation modeling ability for fNIRS data. Our paper demonstrates that the Fuzzy Attention Layer outperforms vanilla self-attention in modeling boundary relations. Additionally, the transformer model with a Fuzzy Attention Layer achieves higher approximation accuracy than the vanilla transformer. This finding is validated with two fNIRS datasets, where relevant neural activity patterns are identified. Given the potential of the Fuzzy Attention Layer, it could be applied to other types of neural data, such as EEG or fMRI, aiding neuroscience and BCI technology in understanding decodable and human-interpretable neural patterns. Our contribution paves the way for significant breakthroughs in understanding the brain’s complex mechanisms and enhancing neurotechnology.
@@ VII Limitations
Acknowledgments: This research was supported by the Institute of Psychology, Chinese Academy of Sciences (Grant No. GJ202007). The research was also sponsored in part by the Australian Research Council (ARC) under discovery grants DP180100656 and DP210101093 and the Australia Defense Innovation Hub under Contract No. P18-650825, US Office of Naval Research Global under Cooperative Agreement Number ONRG - NICOP - N62909-19-1-2058, and AFOSR – DST Australian Autonomy Initiative agreement ID10134. We also thank the NSW Defense Innovation Network and NSW State Government of Australia for financial support in part of this research through grants DINPP2019 S1-03/09 and PP21-22.03.02.

### v2

@@ IV-A3 Participants
In two experiments, fNIRS signals were simultaneously recorded from two participants at a sampling rate of 7.8125 Hz. The device utilized two wavelengths of near-infrared light, 785 nm and 830 nm, with the frequency adjusted according to the wavelength and channel to avoid crosstalk. Each probe was spaced 30 mm apart. For each participant, 8 sources and 7 detectors were used, resulting in 20 source-detector pair channels (CH) per individual, covering the Prefrontal Cortex (PFC). To ensure signal quality, the Scalp Coupling Index (SCI) was calculated for each channel, and those with an SCI below 0.5 were excluded from further analysis. Next, the raw optical signals were converted into relative oxyhaemoglobin (HbO) and deoxyhaemoglobin (HbR) concentration using the Beer-Lambert law, with a partial pathlength factors (ppf) of 0.1. All fNIRS data were then band-pass filtered between 0.01 Hz and 0.2 Hz. Motion artifacts were removed using a rejection criterion as 120 $\mu$mol/L after epoching, and all data were baseline-corrected using a 2-second baseline for future analysis. All preprocessing was performed using the Python MNE package.
@@ IV-B Implementation Details
IV-C Brain Visualization
We utilize MNE and MNE-NIRS[65] in Python to project critical data onto the brain’s surface, enhancing our visualization capabilities. Additionally, to facilitate the identification of frontal lobes in the figures, we delineate the borders of these lobes, as illustrated in all 3D brain figures.
IV-D Evaluation Metrics
@@
As depicted in Table I, we show model performance across two dataset arrangements and the statistical results, each with the channel-first and time-first structures, to investigate the effect of input structuring on the classification outcomes. The dimensional reversal aims to explore the robustness of the model against variations in the temporal-spatial data arrangement and explore the explanation capacity of the Fuzzy Attention Layer.
Given that this study pioneers the classification of handholding conditions using fNIRS data, we establish a baseline with bidirectional LSTM and Transformer models, which are state-of-the-art architectures known for their efficacy in sequence modeling tasks. To validate the efficacy of our proposed approach, we performed an independent t-test to compare it with other models. Each model was tested by training it five times using different random seeds. The Fuzzy Attention Layer-based model achieves the Top 1 in PR AUC metrics in Image Recognition, channel-first, Image Rating channel- and time-first experiments ($p<0.05$).
@@
IV-E Ablation Study
IV-E1 Exploring the Impact of Fuzzy Attention Layer Replacements within Transformer Encoders
@@
IV-E2 Investigating the Rule Count for Model Optimization
The efficacy of the model is analyzed with respect to the number of rules employed within the computational framework. Fig. 6 delineates the relationship between the PR AUC and the rule count, indicating a nuanced dependency on the number of rules. Initially, the PR AUC exhibits an enhancement as the number of rules increases, peaking at approximately ten rules. After that, the model’s performance demonstrates a declining trajectory post the apex, characterized by fluctuations in the PR AUC. This downturn signals a saturation point, beyond which further additions of rules contribute to overfitting. The increase of an excessive number of rules escalates the model’s complexity and parameter count, which, although aimed at refining precision on training data, paradoxically impairs its generalization capabilities on novel, unseen datasets. Consequently, the model exhibits diminished efficacy, underscoring the critical balance required in the rule count to maintain optimal performance without succumbing to overfitting.
@@ IV-E2 Investigating the Rule Count for Model Optimization
IV-E3 Impact of Model Depth on Encoder Performance
@@ V-A Fuzzy Set and Feature Distribution
The fuzzy set distributions of each fuzzy rule, as depicted in Fig. 7, illustrate the temporal dynamics and channel specificity of fuzzy membership functions over multiple experimental conditions. The channels 18, 20, and 29 graphs highlight the variability in attention across different time intervals (3s, 5s, 7s, 10s). Each colored line represents a distinct rule derived from fuzzy logic that interprets the fNIRS data, which is overlaid as blue histograms.
@@
These results show the diversity of the learned fuzzy logic-based rules to dissect the complex temporal patterns.
@@ V-B Sample-wised Interpretability Analysis
To provide an intuitive illustration of how the Fuzzy Attention Layer identifies the handholding condition from fNIRS data, we present the learned center prototype and firing strength for the example fNIRS data in Fig 1C and 1E. To better understand the inference processing, we also present a demo sample from the top-performing model, where the subject holds a hand.
@@ V-C1 Channel-wise Analysis
We selected the model demonstrating the highest performance for the Picture Recognition dataset, which included 10 rules. We chose this top-performing model and used a t-test to analyze the differences in firing strength. The t-values describe the differences between handholding and non-handholding in the PFC, as shown in Fig. 2A. The results indicate how the rules contribute to classification. For example, rules #1, #2, #3, #5, #9, and #10 involve HbO in the right lPFC.
@@ V-C2 Time-wise Analysis
Similarly, in the Picture Recognition dataset, we employed a t-test to evaluate the differences in firing strength between the labels. The t-values illustrating these differences are depicted in Fig. 2B. The results indicate that the crucial time domain is approximately 5 seconds after stimuli presentation in rules #5, #6, #7, and #8, which is highly related to the Haemodynamic Response Function (HRF) [66]. The typical simulated HRF is shown in Fig. 5. This adherence to the HRF is critical as HRF-like data patterns exhibit stable temporal characteristics that are optimally suited for analysis by computational models. Particularly, implementing a fuzzy attention layer within our model shows substantial improvements in performance, as it can more effectively learn and adapt to these HRF-like patterns. Fig. 1C also demonstrates the HRF-like pattern that the center has learned to recognize. This capability underscores the model’s enhanced interpretability and robustness when handling time-first fNIRS data.
Rule #5, in particular, shows the most significant difference between handholding and non-handholding conditions(p < 0.05). Fig. 2C displays the center prototype in PFC for both oxygenated (HbO) and deoxygenated hemoglobin (HbR), and we also observe the left prefrontal cortex (lPFC)’s contribution in the fuzzy centers noted in the Channel-wise model observation.
@@ V-C3 IBS Analysis
The results indicate that both correlation and cosine similarity metrics showed significantly higher IBS in the handholding condition compared to the non-handholding condition ($\textit{p}<0.05$), as shown in Fig. 2E.
@@ V-D Result Explanation
V-E Error Analysis
Figure 9: Error Analysis. A: The percentage of model predictions across different data types using the Fuzzy Attention model and Vanilla Attention model. B: Representative samples for the four possible model prediction conditions ($22x$ 2 grid).
In this section, we analyze the prediction performance of the two best-performing models in our test: the Fuzzy Attention model and the Vanilla Attention model. We evaluate their performance across different data types (emotional image: negative vs. neutral; pair-subject relationship: stranger vs. friend), as shown in Fig. 9A. The results indicate that the Fuzzy Attention model demonstrates a lower percentage of incorrect predictions when processing negative emotional images compared to neutral ones, while it shows minimal variance across different relationship categories. In contrast, the Vanilla Attention model exhibits more uniform performance across both data types.
The superior performance of the Fuzzy Attention model, particularly with negative stimuli, aligns with findings in neuroimaging studies that highlight how negative emotional stimuli elicit distinct neural and physiological responses, especially in regions such as the amygdala and prefrontal cortex, which are responsible for emotion regulation and attention control [76, 77, 78, 79]. This complex interplay between emotion, attention, and physiological responses emphasizes the model’s ability to capture the heightened brain blood response that typically accompanies negative emotional stimuli.
Fig. 9B further analyzes typical samples, comparing the patterns identified by the Fuzzy Attention and Vanilla Attention models. It is evident that when the data aligns closely with HRF, the Fuzzy Attention model successfully predicts the outcome. However, when the data fails to elicit a standard HRF response, possibly due to atypical stimuli, the Fuzzy Attention model struggles to make accurate predictions.
In conclusion, the Fuzzy Attention model excels at recognizing stable neural patterns, particularly those associated with stable responses, such as watching affective stimuli, allowing it to outperform the Vanilla Attention model in scenarios where distinct physiological signals are present.
V-F Future Work
Our proposed model shows a potential fuzzy-rule-based approach to understanding the neural patterns in social neuroscience. While this study has focused on exploring and uncovering neuroscience and psychological theories using fuzzy logic-based models, future work could include optimizing the design of fuzzy centers and automating the decision-making process for the number of fuzzy rules. Additionally, improving the speed of training is essential for enhancing the practical applicability of the model. By addressing these challenges, researchers can enhance the model’s practical applicability and potentially expand its use in real-time social neuroscience studies, paving the way for innovations in therapeutic strategies and human-machine interaction. Future studies should also consider conducting longitudinal research to evaluate the enduring effects of social interactions on neural dynamics, which could further validate and refine the contributions of this research.
@@ VI Conclusion
In this work, we introduce a new attention layer, the Fuzzy Attention Layer, to identify interpretable patterns of neural activity, which are crucial for detailed analysis and understanding of the decision-making process of the neural network for individual cases. This method enhances the correlation modeling ability for fNIRS data. Our paper demonstrates that the Fuzzy Attention Layer outperforms vanilla self-attention in modeling boundary relations. Additionally, the transformer model with a Fuzzy Attention Layer achieves higher approximation accuracy than the vanilla transformer. This finding is validated with two fNIRS datasets, where relevant neural activity patterns are identified. Given the potential of the Fuzzy Attention Layer, its application extends to various neural data types, such as EEG or fMRI, which can significantly enhance neuroscience and BCI technology by decoding and interpreting neural patterns in a human-understandable form. Our contribution sets the stage for substantial advancements in comprehending the brain’s intricate mechanisms and improving neurotechnology. Looking forward, the future scope of this method includes refining the designs of fuzzy centers and the number of rules, accelerating the training process, and implementing the model in real-time social neuroscience settings to explore neuroscience and psychological theories, develop sophisticated therapeutic strategies, and elevate human-machine interactions.